OpenStudy (anonymous):

simplify sqrt(15) * (2sqrt(3)-4sqrt(5))

OpenStudy (anonymous):

=2sqrt(15*3) - 4sqrt(15*5) =2*3sqrt(5) - 4*5sqrt(3) =6sqrt(5)-20sqrt(3)